OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2?

ganeshie8 (ganeshie8):

slope intercept form : \[\large y = \color{Red}{m}x +\color{red}{ b}\] \(\large \color{Red}{m}\) : slope \(\large \color{Red}{b}\) : y intercept

ganeshie8 (ganeshie8):

you're given \(\large \color{red}{m=2} \) : \[\large y = \color{Red}{2}x +\color{red}{ b}\]

ganeshie8 (ganeshie8):

use the given point to find the value of intercept

OpenStudy (anonymous):

3?

ganeshie8 (ganeshie8):

whats 3 ?

OpenStudy (anonymous):

The y intercept.

ganeshie8 (ganeshie8):

nope im getting a different number, how did u get 3 ?

OpenStudy (anonymous):

From the point (3,4).

ganeshie8 (ganeshie8):

you need to plug that point in the equation `y = 2x + b` and solve `b`

ganeshie8 (ganeshie8):

4 = 2*3 + b 4 = 6 + b -2 = b

ganeshie8 (ganeshie8):

so the equation of line would be : \[\large y = \color{Red}{2}x -\color{red}{ 2}\]

OpenStudy (anonymous):

So what you saying is the equation is y = 2x + -2?

ganeshie8 (ganeshie8):

yes but nobody leaves +- like that

ganeshie8 (ganeshie8):

+*- = -

OpenStudy (anonymous):

Alright.